If the square of 2 less than an integer is 16, find the number(s).
Answer by checkley75(3666)

(x-2)^2=16
x^2-4x+4=16
x^2-4x+4-16
x^2-4x-12=0
(x-6)(x+2)=0
x-6=0
x=6 answer.
x+2=0
x=-2 answer.